OpenStudy (anonymous):

| 5x + 13| = –7 show my work

OpenStudy (uri):

You show the work :D

OpenStudy (anonymous):

please help

OpenStudy (uri):

5x =-7 +13 5x =6 x =6/5?

OpenStudy (anonymous):

No solution

OpenStudy (anonymous):

2. Simplify the expression below. 6n2 - 5n2 + 7n2

OpenStudy (uri):

lol dafuq i was wrong then? @Stephensparadox

OpenStudy (uri):

Ask one question at a time @zaz123,Close this ask in a new thread.